OpenStudy (anonymous):

Please Help!!! find the center (h,k) and radius r of circle with given equation. x^2+y^2+10x-16y=-73

OpenStudy (anonymous):

c(-5,-8) and your radius is 4.

OpenStudy (anonymous):

hmmm i don't think that is right

OpenStudy (anonymous):

\[x^2+y^2+10x-16y=-73\] \[x^2+10x+y^2-16y=-73\] now you have to complete the square \[(x+5)^2+(y-8)^2=-73+5^2+8^2=16\] so center is \((-5,8)\) and radius is \(4\)

OpenStudy (rajee_sam):

\[x ^{2} + 10x + y ^{2} - 16y = -73\] Now complete the squares for x and y \[x ^{2} + 10x + 25 + y ^{2} -16y +64 = -73+25+64\] \[(x + 5)^{^{2}} + (y-8)^{2} = 16\] Now this is of the form \[(x-h)^{2} + (y-k)^{2} = r ^{2}\] \[(h,k) \space is \space the \space center \] \[r \space is \space the \space radius\] Center is (-5,8) and radius is 4

OpenStudy (anonymous):

thank you guys so much im not sure who to give the best response to :( you both helped alot.